OpenStudy (blackbird02):

The area of a rectangular lot is increased by 25 m^2 if it is longer by 10 m and narrower by 2.5 m, but when it is shorter by 10m and wider by 5m its area is reduced by 50m^2. Find the area and the dimensions of the lot.

OpenStudy (blackbird02):

How do you solve this?

OpenStudy (anonymous):

(x+10) (y-2.5) = xy + 25 (x-10) (y+5) = xy - 50 solve and x = 20, y = 10

OpenStudy (anonymous):

so area = 20 x 10 = 200

OpenStudy (blackbird02):

@sourwing how did you obtain the values 20 and 10?

OpenStudy (kropot72):

When the two equations of @sourwing are expanded and then simplified, the result is the following two equations: 10y - 2.5x = 50 .........................(1) -10y + 5x = 0 ...........................(2) When equations (1) and (2) are added we get: 2.5x = 50 .........................(3) The solution to equation (3) is: x = 20 m. Plugging this value of x into equation (1) gives y = 10 m.

OpenStudy (blackbird02):

@kropot72 thanks for the explanation

OpenStudy (kropot72):

You're welcome :)